OpenStudy (anonymous):

find the slope and give equation of the line that goes through the points (3,9) and (-2,9)

OpenStudy (anonymous):

slope is 0, correct? what's the next part?

OpenStudy (shamim):

equation of slope of a line is\[m=\frac{ y _{2}-y _{1} }{ x _{2}-x _{1} }\]

OpenStudy (shamim):

here\[x _{1}=3,y _{1}=9,x _{2}=-2,y _{2}=9\]

OpenStudy (anonymous):

So what you do it connect them together

OpenStudy (anonymous):

(9 - 9) / (3 - - 2) = 0 / 5 = 0

OpenStudy (anonymous):

Am I supposed to give the answers on this website, or just hints?

OpenStudy (shamim):

ya

OpenStudy (anonymous):

you can give answer

OpenStudy (anonymous):

0 for slope is what i got...

OpenStudy (anonymous):

same here

OpenStudy (shamim):

and the equation of line will b\[y-y _{1}=\frac{ y _{2}-y _{1} }{ x _{2}-x _{1} }(x-x _{1})\]

OpenStudy (shamim):

just plug in those values in this equation

OpenStudy (anonymous):

plug in the coordinates to that equation?

OpenStudy (shamim):

ya

OpenStudy (shamim):

write the result u got after calculation

OpenStudy (anonymous):

Given points are (3,9) and (-2,9) i.e. x1= 3, x2= -2, y1= 9, y2= 9 \[Slope (m) = \frac{ y _{2} -y _{1}}{ x _{2} -x _{1} }=\frac{ 9 -9}{ -2 -3 }=\frac{ 0}{ -5 }=0\] Now Eq of line passing through two points is given as: \[y-{y _{1}}=\frac{ y _{2}-y _{1} }{ x _{2}-x _{1} }(x-{x _{1}})\rightarrow y-9=0(x-3)\rightarrow y-9=0 \] i.e. y=9 is the required eq of the given points of line.

OpenStudy (anonymous):

ok, that's what i got...so it would be y=9